Which ions are isoelectronic with Xe?

Al3+   K+ F- I- S2- Ba2+
Answer #1

Atomic number of Xe=54

So number of electrons in Xe=54 (as a neutral atom contains same number of electrons as protons as it is electrically neutral)

Now atomic number of Al=13

Charge on a positive ion of an element is due to loss of that many electrons from the neutral atom of that element and charge on a negative ion of an element is due to gain of that many electrons by the neutral atom of that element.

So Al has 13 electrons and Al3+ has 13-3=10 electrons

Atomic number of K=19

K has 19 electrons, K+ has 18 electrons

Atomic number of F=9, F has 9 electrons, so F-​ has 9+1=10 electrons

Atomic number of I=53

I has 53 electrons, I- has 53+1=54 electrons, so I- is isoelectronic with Xe.

Atomic number of S=16

S has 16 electrons, S2- has 16+2=18 electrons

Atomic number of Ba=56

Number of electrons in Ba=56, and number of electrons in Ba2+=56-2=54 electrons. So Ba2+ is isoelectronic with Xe.

Thus both I- and Ba2+ are isoelectronic with Xe